OpenStudy (anonymous):

x=5y-8 This is the inverse of f(x) = 5x - 8, correct? And I need help solving for y :)

OpenStudy (anonymous):

No.. To get the inverse of f(x) = 5x - 8, solve for x first y = 5x - 8 5x = y + 8 x = (y + 8) /5 Now switch the y and x y = (x + 8) /5 f(x) = (x + 8) /5 is your inverse

OpenStudy (johnweldon1993):

Well yes that is the first step...switching the positions of 'x' and 'y'... so now you have x = 5y - 8 And you want to solve for 'y' again, how would you do so?

OpenStudy (anonymous):

I'm not sure...Math has never been my strength in school. :/

OpenStudy (anonymous):

I'm a good learner though @johnweldon1993

OpenStudy (johnweldon1993):

Well you can do either this first or as BangKokGarret has posted above If you do this the way I have...you start with x = 5y - 8 Now use algebra to get the 'y' by itself...add 8 to both sides to isolate the 5y first x = 5y - 8 +8 +8 ----------- x + 8 = 5y Now you just divide everything by 5 to completely solve for 'y' x + 8 = 5y ---------- 5 5 y = x+8 ---- 5

OpenStudy (anonymous):

So you got that step by dividing? Hence the fraction?

OpenStudy (anonymous):

@johnweldon1993 thanks, I think I'm good to go. Appreciate the help.

OpenStudy (johnweldon1993):

Yeah dividing...just because as you see the 'y' is being multiplied by 5...so to cancel that out...you DIVIDE by 5